Setting locant numbers by locant.sty (v1.01) of XΥMTEX

Shinsaku Fujita Ashigara Research Laboratories, Fuji Photo Film Co., Ltd., Minami-Ashigara, Kanagawa, 250-01 Japan

1996/08/16

1 Introduction

1.1 Options for docstrip

option function
locant locant.sty
driver driver for this dtx file

1.2 Version Information

1 ⟨∗locant⟩

2 \typeout{XyMTeX for Drawing Chemical Structural Formulas. Version 1.01}
3 \typeout{       -- Released August 16, 1996 by Shinsaku Fujita}
4 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
5 \def\j@urnalname{locant}
6 \def\versi@ndate{August 16, 1996}
7 \def\versi@nno{ver1.01}
8 \def\copyrighth@lder{SF} % Shinsaku Fujita
9 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
10 \typeout{XyMTeX Macro File `\j@urnalname' (\versi@nno) <\versi@ndate>%
11 \space[\copyrighth@lder]}

2 Macros for designating locant numbers

2.1 Bond indicators for vertical formulas

\dbloocant \dblocant The macro \dbloocant is used to designate six edges (inner double bonds) of a six-membered ring. Any character can be asigned to the respective bond. The macro \dblocant provides a fixed set of aliphabetical charactors on the six edges.

12 \def\bdloocant#1#2#3#4#5#6{%
13 \begin{picture}(800,880)(-\shiftii,-\shifti)
14 \putratom{90}{360}{#1}%

15 \putratom{200}{180}{#2}%
16 \putratom{90}{-20}{#3}%
17 \putlatom{-90}{360}{#6}%
18 \putlatom{-200}{180}{#5}%
19 \putlatom{-90}{-20}{#4}%
20 \end{picture}}%
21 \def\bdlocant{\bdloocant{a}{b}{c}{d}{e}{f}}

\sxloocant \sxlocant The macro \sxloocant is used to designate six vertices of a six-membered ring. Any character can be asigned to the respective bond. The macro \sxlocant provides a fixed set of arabic numerals on the six edges.

22 \def\sxloocant#1#2#3#4#5#6{%
23 \begin{picture}(800,880)(-\shiftii,-\shifti)%
24 \putratom{-10}{340}{\scriptsize #1}%
25 \putlatom{150}{260}{\scriptsize #2}%
26 \putlatom{150}{110}{\scriptsize #3}%
27 \putratom{-10}{40}{\scriptsize #4}%
28 \putratom{-150}{110}{\scriptsize #5}%
29 \putratom{-150}{260}{\scriptsize #6}%
30 \end{picture}}%
31 \def\sxlocant{\sxloocant{1}{2}{3}{4}{5}{6}}

2.2 Bond indicators for vertical formulas

\dbloocanth \dblocanth The macro \dbloocanth is used to designate six edges (inner double bonds) of a horizontal six-membered ring. Any character can be asigned to the respective bond. The macro \dblocanth provides a fixed set of aliphabetical charactors on the six edges.

32 \def\bdloocnth#1#2#3#4#5#6{%
33 \begin{picture}(880,800)(-\shifti,-\shiftii)%
34 \putlatom{10}{80}{#1}%
35 \putratom{180}{210}{#2}%
36 \putratom{380}{80}{#3}%
37 \putratom{380}{-120}{#4}%
38 \putratom{180}{-260}{#5}%
39 \putlatom{10}{-120}{#6}%
40 \end{picture}}%
41 \def\bdlocnth{\bdloocnth{a}{b}{c}{d}{e}{f}}

\sxloocanth \sxlocanth The macro \sxloocanth is used to designate six vertices of a horizontal sixmembered ring. Any character can be asigned to the respective bond. The macro \sxlocanth provides a fixed set of arabic numerals on the six edges.

42 \def\sxloocnth#1#2#3#4#5#6{%
43 \begin{picture}(880,800)(-\shifti,-\shiftii)%
44 \putratom{40}{-20}{\scriptsize #1}%
45 \putratom{110}{100}{\scriptsize #2}%
46 \putlatom{290}{100}{\scriptsize #3}%
47 \putlatom{380}{-20}{\scriptsize #4}%

48 \putratom{110}{-140}{\scriptsize #6}%
49 \putlatom{290}{-140}{\scriptsize #5}%
50 \end{picture}}%
51 \def\sxlocnth{\sxloocnth{1}{2}{3}{4}{5}{6}}

52 ⟨/locant⟩